# Frequency-Assisted Mamba for Remote Sensing Image Super-Resolution

### v1

@@ IV-D1 Effect of Key Modules
(a) Effect of VSSM. Table I reports that the Model-1 (baseline) obtains 27.751 dB. Model-2 exhibits a performance gain of 0.089 dB over the baseline model, which demonstrates the effectiveness of VSSM in global modeling. In addition, in Fig. 5, we visualize the intermediate feature maps of the baseline model and our FMSR. The features are obtained by visualizing the 56th channels at the end of FMG. With VSSM, the results of our FMSR are more prominent across the entire feature maps, highlighting the favorable global modeling capability.
(b) Effect of HGM. Employing both VSSM and the MLP layer, Model-3 demonstrates a favorable performance improvement of 0.258 dB. However, when equipped with the channel attention mechanism to introduce more locality, Model-4 performs similarly to Model-2 (28.088 dB vs. 28.104 dB). This suggests that neither MLP nor channel attention is less effective in boosting the performance of Mamba. In this context, further improving the reconstruction performance becomes very challenging. Our HGM achieves a performance gain of 0.034 dB. Thus, to benefit from both global and local inductive bias, we combine VSSM and HGM in our FMSR.

### v2

@@ IV-D1 Effect of Key Modules
(a) Effect of VSSM. Table I reports that the Model-1 (baseline) obtains 27.751 dB. Model-2 exhibits a performance gain of 0.089 dB over the baseline model, which demonstrates the effectiveness of VSSM in global modeling. To demonstrate the linear complexity of VSSM, we conducted complexity experiments with inputs of different resolutions, as shown in Fig. 8. Specifically, we adopted the standard Multihead Self-Attention (MSA) [71] with a dimension of 180 as the baseline and adjusted our model to have a dimension of 144. This adjustment ensured that the complexity in terms of parameters (0.1308M vs. 0.1279M) and FLOPs (0.1534G vs. 0.1312G) was roughly equivalent. Our method is more efficient than the widely used MSA and exhibits linear complexity with input resolution. In addition, in Fig. 5, we visualize the intermediate feature maps of the baseline model and our FMSR. The features are obtained by visualizing the 56th channels at the end of FMG. With VSSM, the results of our FMSR are more prominent across the entire feature maps, highlighting the favorable global modeling capability.
(b) Effect of HGM. Employing both VSSM and the MLP layer, Model-3 demonstrates a favorable performance improvement of 0.258 dB. However, when equipped with the channel attention mechanism to introduce more locality, Model-4 performs similarly to Model-2 (28.088 dB vs. 28.104 dB). This suggests that neither MLP nor channel attention is less effective in boosting the performance of Mamba. In this context, further improving the reconstruction performance becomes very challenging. Our HGM achieves a performance gain of 0.034 dB. Thus, to benefit from both global and local inductive bias, we combine VSSM and HGM in our FMSR. In addition, we have investigated the impact of different expansion factors $\lambda$ used in the linear projection layer. The quantitative results are presented in Fig. 7. It is observed that the performance of FMSR does not fluctuate significantly with changes in $\lambda$. We ultimately selected $\lambda=2$ as it delivers the best performance.
@@ V Conclusion
As prior research demonstrated that achieving optimal performance on the ×$\times 4$ SR task typically translates to favorable performance on lower scaling factors, this study focused solely on the ×$\times 4$ SR. While saving computational resources, this specificity lacks flexibility in exploring SR at different scales. In future work, we plan to extend our FMSR to include more scaling factors, further demonstrating its robustness and effectiveness.